OpenStudy (amistre64):

Before a market study, there is a 70% chance that the market will be good. When the market is good, the study has a track record of predicting a good market 85% of the time. When the market is bad, the study has a track record of predicting a good market 20% of the time. Given that the study has predicted a good market, what is the probability that the market is actually good?

OpenStudy (amistre64):

or rather: .....what is the probability that the market will actually be good?

OpenStudy (anonymous):

baye's formula?

OpenStudy (amistre64):

yep

OpenStudy (amistre64):

i tend to do a table, rather than the formula itself ...

OpenStudy (amistre64):

i get lost in the P(B|A')s and P(A|B)s and such lol

OpenStudy (anonymous):

actually that is always confusing. i am going to write down what i think the answer is, and then how i got it i think it is \[\frac{.85\times .7}{.85\times .7+.2\times .3}\]

OpenStudy (amistre64):

\begin{array}l &&&&prior:&70\%&30\%\\ &+&-&\\ good\ markt&[85\%]&15\%&\\ bad\ market&20\%&80\%&\\ \end{array} .85*.70 = .5950 / X <--- +.20*.30 = .0600 ---------------- X = .6550 P(answer) = .595/.060 yep, same answers :)

OpenStudy (anonymous):